# ChangeMamba: Remote Sensing Change Detection with Spatio-Temporal State Space Model

### v1

@@ Abstract
Convolutional neural networks (CNN) and Transformers have made impressive progress in the field of remote sensing change detection (CD). However, both architectures have their inherent shortcomings. Recently, the Mamba architecture, based on spatial state models, has shown remarkable performance in a series of natural language processing tasks, which can effectively compensate for the shortcomings of the above two architectures. In this paper, we explore for the first time the potential of the Mamba architecture for remote sensing change detection tasks. We tailor the corresponding frameworks, called MambaBCD, MambaSCD, and MambaBDA, for binary change detection (BCD), semantic change detection (SCD), and building damage assessment (BDA), respectively. All three frameworks adopt the cutting-edge visual Mamba architecture as the encoder, which allows full learning of global spatial contextual information from the input images. For the change decoder, which is available in all three architectures, we propose three spatio-temporal relationship modeling mechanisms, which can be naturally combined with the Mamba architecture and fully utilize its attribute to achieve spatio-temporal interaction of multi-temporal features and obtain accurate change information. On five benchmark datasets, our proposed frameworks outperform current CNN- and Transformer-based approaches without using any complex strategies or tricks, fully demonstrating the potential of the Mamba architecture. Specifically, we obtained 83.11, 88.39 and 94.19 F1 scores on the three BCD datasets SYSU, LEVIR-CD+, and WHU-CD; on the SCD dataset SECOND, we obtained 24.04 SeK; and on the xBD dataset, we obtained 81.41 overall F1 score. The source code will be available in https://github.com/ChenHongruixuan/MambaCD
@@
Change detection (CD) has been a popular field within the remote sensing community since the inception of remote sensing technology. It aims to detect changes in objects on the Earth’s surface from multi-temporal remote sensing images acquired at different times. Depending on the desired result from the change detector, CD tasks can be categorized into three types, namely binary CD (BCD), semantic CD (SCD), and building damage assessment (BDA). Now, CD techniques play an important role in many fields, including land cover change analysis, urban sprawl studies, disaster response, geographic information system (GIS) updating and ecological monitoring [1, 2, 3, 4, 5, 6].
@@
Optical high-resolution remote sensing imagery has become one of the most applied and researched data sources in the field of CD, which can provide detailed textural and geometric structural information about surface features, allowing us to achieve more refined changes. Due to the increased heterogeneity within the same land-cover feature brought about by the increase in spatial resolution, the traditional pixel-based CD methods [7, 8, 9] are difficult to achieve satisfactory detection results. To cope with this, researchers proposed object-based CD methods [10], which take the object consisting of homogeneous pixels as the basic unit for CD. However, these methods rely on hand-designed “shallow” features, which are not representative and robust enough to cope with complex ground conditions in multi-temporal high-resolution images [11].
@@
The emergence of deep learning has brought new models and paradigms for CD, greatly improving the efficiency and accuracy of CD. After the early days of developing image patch-based approaches with tiny deep learning models [12, 13, 14], the released of a number of large-scale benchmark datasets [15] allowed us to start designing and introducing even larger and deeper models. Convolutional neural network (CNN)-based methods have been dominant within the field of CD after the introduction of fully convolutional networks (FCNs) to the field by Daudt et al. [16]. By employing models within the computer vision field as well as a priori knowledge of the CD task, many representative methods have been proposed [17, 18, 3, 4, 19]. Despite the decent results achieved by these methods, the inherent shortcomings of the CNN model structure, i.e., the inability of the limited receptive field to capture long-range dependencies between pixels, make these methods still fall short when dealing with complex and diverse multi-temporal scenes in images with different spatial-temporal resolutions [20]. The emergence of visual Transformer [21] provides an effective way to solve the above problems. By means of the self-attention operation, the transformer architecture can efficiently model the relationship between all the pixels of the whole image. Currently, more and more CD architectures adopt Transformer architectures as the encoder to extract representative and robust features [22, 23, 24] and utilize it in the decoder to capture the spatio-temporal relationship between multi-temporal features [20, 25], exhibiting superior performance.
@@
Nonetheless, the self-attention operation requires quadratic complexity in terms of image sizes, resulting in expensive computational overhead, which is detrimental in dense prediction tasks, e.g., land-cover mapping, object detection, and CD, in large-scale remote sensing datasets. Some available solutions, such as limiting the size of the computational window or step size to improve the efficiency of attention, are at the cost of imposing limits on the size of the reception field [26, 27]. As a viable alternative to the Transformer architecture, state space sequence models (SSMs), especially structured state space sequence models (S4) [28], have shown cutting-edge performance in continuous long-sequence data analysis, with the promising attribute of linearly scaling in sequence length. The Mamba architecture [29] further improves on S4 by utilizing a selection mechanism that allows the model to select relevant information in an input-dependent manner. By combining it with hardware-aware implementations, Mamba outperforms Transformers on a number of downward exploration tasks with high efficiency. Very recently, Mamba architecture has been extended to image data and shown promising results on some visual tasks [30, 31]. However, its potential for high-resolution optical images and different CD subtasks is still under-explored. How to design a suitable architecture based on a priori knowledge of CD tasks is undoubtedly valuable for subsequent research in the community.
@@
In this paper, we present the ChangeMamba architecture, an efficient Mamba model for remote sensing CD tasks. ChangeMamba can efficiently model the global spatial context, thereby showing very effective results on the three subtasks of CD, namely BCD, MCD, and BDA. Specifically, ChangeMamba is based on the recently proposed VMamba architecture [30], which adopts a Cross-Scan Module (CSM) to unfold image patches in different spatial directions to achieve effective modeling of global contextual information of images. In addition, since CD tasks require the detector to adequately capture the spatio-temporal relationships between multi-temporal images, we design a cross-spatio-temporal state space module (CST-Mamba), which can adequately model the spatio-temporal relationships between multi-temporal features by combining them in different ways, and thus efficiently detect the different categories of changes, including binary changes, semantic changes, and building damage levels.
@@
For the first time, we explore the application of the Mamba architecture in the field of remote sensing change detection, thereby achieving high accuracy and efficient CD.
@@
Based on the Mamba architecture and incorporating the unique characteristics of CD, we design the corresponding network frameworks for each of the three CD tasks: BCD, SCD, and BDA.
@@ II-A CNN-Based Method
The rapid evolution of CNN and their prowess in extracting local features have led to early applications in CD tasks. In the realm of CNN-based BCD, an initial and simple approach [16] involved combining bi-temporal images into a single input for processing by an FCN. In addition, it introduced two variations that employ siamese encoders to handle dual inputs. Given the roots of FCNs in semantic segmentation, [17] sought to tailor these networks more closely to BCD tasks by integrating feature change maps as a deep supervised signal, thus refining the final prediction. Addressing the limited receptive field of CNNs, [13] expanded this framework by combining long- and short-term memory (LSTM) networks with CNNs. Furthermore, Fang et al. [32] designed a Siamese architecture with denser connections to facilitate more effective shallow information exchange for BCD. To combat the challenge of uneven foreground-background distributions within BCD, Han et al. [33] employed a novel sampling approach and an attention mechanism to integrate global information more effectively. After that, they further introduced an innovative self-attention module that not only enhances long-distance dependency, but also specifically targets the refinement of BCD predictions, significantly improving the delineation of change area boundaries and reducing the occurrence of inner holes [34].
@@
Compared to BCD, which aims to identify ”where” changes have occurred, SCD seeks to determine ”what” the changes are. Although this poses a greater challenge, it is more important for downstream applications in remote sensing. In the field of CNN-based SCD, FCNs remain the core structure. Early explorations in SCD include [35], which adopts a multitask learning framework to simultaneously predict a binary change mask and land cover mapping, using the latter to assist in generating the final semantic change prediction. The approach in [36] represents an early attempt to combine recurrent neural network (RNN) and CNN with SCD, where CNN extracts semantic features and RNN models temporal dependencies to classify multiple types of changes. In [37] and [11], unsupervised SCD methods were developed, using pre-trained CNNs or unsupervised learning techniques to extract temporal features, then comparing these features to classify changes. [38] introduced a benchmark dataset for SCD along with a new evaluation metric. Building on [35], Ding et al. [39] proposed a more innovative architecture to address the insufficient exchange of information between the semantic and change branches, introducing new loss functions and supervision signals to force the network to correlate two branches. Zheng et al. [18] continues to strengthen the connection between the two branches, designing a temporal-symmetric transformer (TST) module to model the causal relationship between semantic and change representations. In [40], channel attention is used to embed change information into temporal features. In [41], temporal salient features are used and fused to extract and improve the representation of changes.
@@
Compared to the advances in CNN-based methods for BCD and SCD, the progression of BDA methodologies has been relatively gradual. BDA intricately categorizes areas affected by disasters into four distinct classes: ”No damage”, ”Minor damage”, ”Major damage”, and ”Destroyed”. The procedural approach to BDA typically encompasses two primary phases: identifying the location of the building and assessing the level of damage incurred. This bifurcation essentially positions BDA as a nuanced subset of SCD, with a focus on quantifying damage levels. Among established approaches, the xView2 challenge [42] introduced a baseline method that employs a UNet+ResNet cascade for building localization and damage classification, trained with pre and post-disaster imagery. The Siamese-UNet, the winning approach of the xView2 challenge [42], utilizes a UNet framework for both pixel-level building localization and damage level classification. It innovatively initializes the damage classification UNet with the parameters from the building localization UNet, enhancing performance through a multi-model ensemble of four UNets. Further advancements [43] draw from SCD solutions and the Mask-RCNN architecture to address BDA, incorporating Siamese encoders, multitask heads, and instance-level predictions. An evolved approach [3] introduces a hybrid encoder architecture that merges shared and task-specific weights, alongside an FCN-based multitask decoder. This development enhances task integration with an object-based post-processing strategy, resulting in more refined outcomes.
@@
The development of CNN-based CD methods has matured significantly over time. However, despite their excellent ability to extract local features, CNN architectures inherently struggle to capture long-distance dependencies. This limitation becomes particularly critical in CD tasks, where change regions, compared to background areas, are often sparse and scattered. The network’s ability to model relationships beyond local areas is essential for accurately identifying these change regions. In this context, Mamba, evolved from RNN architectures, stands out for its exceptional ability to model non-local relationships, making it ideally suited for CD tasks. Its architecture leverages the sequential processing strength of RNNs, enabling it to effectively understand the spatial context and dependencies over larger areas, thus overcoming one of the key limitations of traditional CNN-based approaches in CD.
@@ II-B Transformer-Based Method
Recently, with the rise of transformers in computer vision tasks, their superior long-distance modeling capabilities, compared to CNNs, have made them particularly effective in the field of BCD, SCD, and BDA. In the realm of BCD, [20] emerges as a pioneering approach, transforming images into semantic tokens to model spatial-temporal contexts within a token-based framework, enhancing the change detection process. Bandara et al. [22] employing a pure transformer-based Siamese network that uses a structured transformer encoder coupled with an MLP decoder. This architecture eliminates the need for a CNN-based feature extractor. At the same time, Zhang et al. [24] also proposes a pure transformer-driven model, adopting a Siamese U-shaped structure, processing images in parallel to extract and merge multiscale features. It leverages Swin transformer blocks across its architecture, further demonstrating the transformer’s capability to refine change detection outcomes. Li et al. [25] introduces a hybrid model that blends understanding of the global context of transformers with the semantic detail resolution of UNet. It eliminates UNet’s redundant information and enhances feature quality through a difference enhancement module, leading to more precise change area delineation.
@@
As for transformer-based SCD, Niu et al. [44] uses a PRTB backbone, combining preactivated residual and transformer blocks, to extract semantic features from image pairs, which is crucial for detecting subtle changes. Its Multi-Content Fusion Module (MCFM) precisely differentiates between the foreground and background. The model’s efficacy is further amplified by multitask prediction branches and a custom loss function, ensuring nuanced semantic change detection with heightened accuracy. Alternatively, Ding et al. [45] introduces the Semantic Change Transformer (SCanFormer), a CSWin Transformer adaptation, to explicitly chart semantic transitions. The model utilizes a triple encoder-decoder architecture to improve spatial features, and it is distinguished by incorporating spatio-temporal dynamics and task-specific expertise, leading to a decrease in learning disparities. This approach significantly boosts accuracy on benchmarks, adeptly managing scenarios with sparse change samples and semantic labels.
@@
In response to the transformer-based BDA, Chen et al. [23] uses a siamese Mix Transformer (MiT) encoder and lightweight All-MLP decoder efficiently to process bi-temporal image pairs. It features deep feature extraction and a multitemporal fusion module for improved information integration, and a dual-task decoder for precise building localization and damage classification. By adopting strategies like weight sharing and innovative fusion mechanisms, it efficiently handles high-resolution imagery with lower computational costs, making it effective and streamlined for BDA.
@@
Although transformer-based methods have significantly enhanced CD accuracy through their powerful non-local modeling capabilities, they come with substantial computational resource demands. Specifically, as the number of tokens increases, the consumption of computational resources escalates quadratically, posing challenges for processing large-scale input. However, Mamba, an architecture inherited from RNNs, not only rivals transformers in non-local modeling capabilities but also maintains computational resource consumption at a linear growth rate.
Moreover, spatio-temporal fusion remains a fundamental aspect of CD tasks, where seamless and efficient integration of time and space information is essential to achieve accurate detection results. Traditional methods, whether they employ CNNs or transformers, tend to follow a rigid, modular stacking approach for spatio-temporal fusion. In contrast, the Mamba framework introduces a novel paradigm by treating tokens with inherent ”sequence,” offering a dynamic range of spatio-temporal fusion possibilities. By manipulating the order in which tokens are processed, Mamba can generate a spectrum of fusion outcomes, marking a significant departure from conventional methods. This innovative approach to spatiotemporal fusion provides Mamba with unparalleled flexibility and effectiveness in addressing CD tasks.

### v2

@@ Abstract
Convolutional neural networks (CNN) and Transformers have made impressive progress in the field of remote sensing change detection (CD). However, both architectures have inherent shortcomings. Recently, the Mamba architecture, based on state space models, has shown remarkable performance in a series of natural language processing tasks, which can effectively compensate for the shortcomings of the above two architectures. In this paper, we explore for the first time the potential of the Mamba architecture for remote sensing CD tasks. We tailor the corresponding frameworks, called MambaBCD, MambaSCD, and MambaBDA, for binary change detection (BCD), semantic change detection (SCD), and building damage assessment (BDA), respectively. All three frameworks adopt the cutting-edge Visual Mamba architecture as the encoder, which allows full learning of global spatial contextual information from the input images. For the change decoder, which is available in all three architectures, we propose three spatio-temporal relationship modeling mechanisms, which can be naturally combined with the Mamba architecture and fully utilize its attribute to achieve spatio-temporal interaction of multi-temporal features, thereby obtaining accurate change information. On five benchmark datasets, our proposed frameworks outperform current CNN- and Transformer-based approaches without using any complex training strategies or tricks, fully demonstrating the potential of the Mamba architecture in CD tasks. Specifically, we obtained 83.11, 88.39 and 94.19 F1 scores on the three BCD datasets SYSU, LEVIR-CD+, and WHU-CD; on the SCD dataset SECOND, we obtained 24.11 SeK; and on the BDA dataset xBD, we obtained 81.41 overall F1 score. Further experiments show that our architecture is quite robust to degraded data. The source code is available in https://github.com/ChenHongruixuan/MambaCD.
@@
Change detection (CD) has been a popular field within the remote sensing community since the inception of remote sensing technology. It aims to detect changes in objects on the Earth’s surface from multi-temporal remote sensing images acquired at different times. Depending on the desired result from the change detector, CD tasks can be categorized into three types, namely binary CD (BCD), semantic CD (SCD), and building damage assessment (BDA). Now, CD techniques play an important role in many fields, including land cover change analysis, urban sprawl studies, disaster response, geographic information system (GIS) updating and ecological monitoring [1, 2, 3, 4, 5, 6, 7].
@@
Optical high-resolution remote sensing imagery has become one of the most applied and researched data sources in the field of CD, which can provide detailed textural and geometric structural information about surface features, allowing us to achieve more refined changes [3]. Due to the increased heterogeneity within the same land-cover feature brought about by the increase in spatial resolution, traditional pixel-based CD methods [8, 9, 10] are difficult to achieve satisfactory detection results. To cope with this, researchers proposed object-based CD methods [11, 12, 13], which take the object consisting of homogeneous pixels as the basic unit for CD. However, these methods rely on hand-designed “shallow” features, which are not representative and robust enough to cope with complex ground conditions in multi-temporal high-resolution images [14, 15].
@@
The emergence of deep learning has brought new models and paradigms for CD, greatly improving the efficiency and accuracy of CD. After the early days of developing image patch-based approaches with tiny deep learning models [16, 17, 18], the released of a number of large-scale benchmark datasets [19, 20, 21] have allowed us to start designing and introducing larger and deeper models. Convolutional neural network (CNN)-based methods have been dominant within the field of CD since the introduction of fully convolutional networks (FCNs) to the field by Daudt et al. [22]. By employing models within the computer vision field as well as a priori knowledge of the CD task, many representative methods have been proposed [23, 24, 3, 4, 25, 26, 27]. Despite the decent results achieved by these methods, the inherent shortcomings of the CNN model structure, i.e., the inability of the limited receptive field to capture long-range dependencies between pixels, make these methods still fall short when dealing with complex and diverse multi-temporal scenes in images with different spatial-temporal resolutions [28]. The emergence of visual Transformer [29] provides an effective way to solve the above problems. By means of the stacked self-attention modules, the Transformer architecture can efficiently model the relationship between all the pixels of the whole image. Currently, more and more CD architectures adopt Transformer architectures as the encoder to extract representative and robust features [30, 31, 32] and utilize it in the decoder to capture the spatio-temporal relationships between multi-temporal features [28, 33], exhibiting superior performance.
@@
Nonetheless, the self-attention operation demands quadratic complexity in terms of image sizes, resulting in expensive computational overhead, which is detrimental in dense prediction tasks, e.g., land-cover mapping, object detection, and CD, in large-scale remote sensing datasets. Some available solutions, such as limiting the size of the computational window or step size to improve the efficiency of attention, are at the cost of imposing limits on the size of the reception field [34, 35]. As a viable alternative to the Transformer architecture, state space sequence models (SSMs), especially structured state space sequence models (S4) [36], have shown cutting-edge performance in continuous long-sequence data analysis, with the promising attribute of linearly scaling in sequence length. The Mamba architecture [37] further improves on S4 by employing a selection mechanism that allows the model to select relevant information in an input-dependent manner. By combining it with hardware-aware implementations, Mamba outperforms Transformers on a number of downstream tasks with high efficiency. Very recently, Mamba architecture has been extended to image data and shown promising results on some visual tasks [38, 39]. However, its potential for high-resolution optical images and different CD subtasks is still under-explored. How to design a suitable architecture based on a priori knowledge of CD tasks is undoubtedly valuable for subsequent research in the community.
@@
In this paper, we present the first attempt at introducing Mamba architecture for change detection tasks and proposed several spatio-temporal state space models (collectively called ChangeMamba) to realize effective and robust CD. ChangeMamba can effectively model the global spatial context and spatio-temporal relationships in multitemporal images, thereby showing very promising results on the three subtasks of CD, namely BCD, MCD, and BDA. Specifically, ChangeMamba is based on the recently proposed VMamba architecture [38], which adopts a Cross-Scan Module (CSM) to unfold image patches in different spatial directions to achieve effective modeling of global contextual information of images. Since CD tasks require the detector to adequately learn spatio-temporal features from multi-temporal images, we design three spatio-tepmoral relationship modelling mechanisms and naturally combine it with Mamba architecture, which can adequately model the spatio-temporal relationships between multi-temporal features by scanning them in different ways, and thus efficiently detect the different categories of changes, including binary changes, semantic changes, and building damage levels.
@@
For the first time, we explore the application of the Mamba architecture in the field of remote sensing change detection, thereby achieving high accuracy, efficient, and robust CD.
@@
Based on the Mamba architecture and incorporating the unique characteristics of CD tasks, we design the corresponding network frameworks for each of the three CD tasks: BCD, SCD, and BDA.
@@ II-A CNN-Based Method
The rapid evolution of CNN and their prowess in automatically extracting hierarchical features have led to wide applications in CD tasks. After the early stage of patch-based framework [16, 18], Daudt et al. [22] presented the first attempt at introducing FCNs into the BCD field and three basic FCNs architecture were developed. After that, many representative network architectures were proposed. Zhang et al. [23] sought to tailor the networks for semantic segmentation task more suitable for the BCD task by integrating feature change maps as a deep supervised signal, thus refining the final prediction. To make the framework suitable for handling multi-source images, Chen et al. [17] proposed a framework by combining long- and short-term memory (LSTM) networks with CNNs.Furthermore, Fang et al. [40] designed a Siamese architecture with denser connections to facilitate more effective shallow information exchange for BCD. To combat the challenge of uneven foreground-background distributions within BCD, Han et al. [41] employed a novel sampling approach and an attention mechanism to integrate global information more effectively. After that, they introduced an innovative self-attention module that not only enhances long-distance dependency, but also specifically targets the refinement of BCD predictions, significantly improving the delineation of change area boundaries and reducing the occurrence of inner holes [42]. Zhao et al. [43] explored the effect of different fusion strategies on BCD performance and designed a dual encoder-decoder architecture to enhance the detection performance. To solve the problem of insufficient sample, the transfer learning technique was also introduced. Guo et al. [4] developed an attention-based network for semi-supervised building updating. Cao et al. [25] designed a full-level fused cross-task transfer learning architecture for building change detection. Recently, Chen et al. [44] has developed a sample generation method based on simply exchanging image patches, which can train change detectors on single-temporal images in unsupervised manner.
@@
Compared to BCD, which only aims to identify “where” changes have occurred, SCD further seeks to determine “what” the changes are. Although this poses a greater challenge, it is more important for practical application analysis [45, 24]. Early explorations based on CNNs for SCD is the study [45], which adopts a multitask learning framework to simultaneously predict binary change maps and land cover maps, using the latter to assist in generating the final semantic change prediction. Mou et al. [46] proposed to combine recurrent neural network (RNN) and CNN for SCD, where a Siamese CNN extracts semantic features and RNN models temporal dependencies to classify multiple types of changes. Yang et al. [20] introduced a benchmark dataset for SCD along with a new evaluation metric. Based on the framework in [45], Ding et al. [47] proposed a more powerful architecture to address the insufficient exchange of information between the semantic and change decoders. A new loss functions was introduced to force the network to correlate two decoders. Zheng et al. [24] continues to strengthen the interactions between multi-temporal features, designing a temporal-symmetric transformer (TST) module to model the spatio-temporal relationships. In [48], channel attention was used to embed change information into temporal features. In [49], temporal salient features were fused to extract and improve the representation of changes. In [50] and [14], unsupervised SCD methods were also studied, using pre-trained CNNs or unsupervised learning techniques to extract hierarchical features, then comparing these features to determine changes.
@@
Due to lack of suitable benchmark datasets caused by the scarcity of disaster events and difficulty in annotation, the methodology progression of BDA has been relatively slow. The establishment of xBD benchmark dataset of the xView2 Challenge [51], in turn, provides a large-scale, common benchmark for the field. The procedural approach to BDA typically encompasses two primary phases: identifying the location of the building and assessing the level of damage incurred. In the xBD dataset, there are four damahge levels, i.e., “no damage”, “minor damage”, “major damage”, and “destroyed”. This bifurcation essentially positions BDA as a nuanced subset of SCD, with a focus on quantifying damage levels. Among established approaches, the xView2 challenge [51] introduced a baseline method that employs a residual UNet for building localization and damage classification, trained with pre and post-disaster imagery. The Siamese-UNet, the winning approach of the xView2 challenge [51], utilized a UNet framework for both pixel-level building localization and damage level classification. It initialized the UNet for the damage classification task with the parameters from the UNet for the building localization task. Further advancements [52] utilized Mask-RCNN architecture [53] to address BDA, incorporating Siamese encoders, multitask heads, and instance-level predictions. Zheng et al. [3] introduced a deep object-based architecture that encompasses shared and task-specific encoders, alongside an FCN-based multitask decoder. This development enhances task integration with an object-based post-processing strategy, resulting in more refined outcomes.
@@
The development of CNN-based CD methods has matured significantly over time. However, despite their excellent ability to extract local features, CNN architectures inherently struggle to capture long-distance dependencies [29]. This limitation becomes particularly critical in CD tasks, where change regions, compared to background areas, are often sparse and scattered. The network’s ability to model relationships beyond local areas is essential for accurately identifying these changed regions [28, 30].
@@ II-B Transformer-Based Method
Recently, with the promising results obtained by Transformers in the computer vision fields [29, 35], their superior long-distance modeling capabilities, have made them effectively overcome the challenges faced by the CNNs. As a result, many studies have begun to address CD tasks using Transformer architectures. In the field of BCD, Chen et al. [28] emerged as a pioneering approach, transforming multi-temporal images into semantic tokens to model spatial-temporal relationships within a token-based framework, enhancing the change detection process. Bandara et al. [30] proposed a pure transformer-based Siamese network, which has a Siamese Transformer encoder coupled with an multilayer perceptron (MLP) decoder. This architecture eliminates the need for a CNN-based feature extractor. Similarly, Zhang et al. [32] proposed a pure Transformer-driven model, adopting a Siamese U-shaped structure to fully recover detailed change information. Specifically, it leveraged the Swin Transformer blocks [34] across its architecture, further demonstrating the transformer’s capability in BCD tasks. Li et al. [33] introduced a hybrid model that blends the global contextual modelling capability of Transformers with the CNN block and UNet architecture. It eliminates UNet’s redundant information and enhances feature quality through a difference enhancement module, leading to more precise detection results.
@@
For Transformer-based SCD, Niu et al. [54] used a PRTB backbone, combining preactivated residual and Transformer blocks, to extract semantic features from multi-temporal image pairs, which is crucial for detecting subtle changes. The model’s efficacy is further amplified by multitask prediction branches and a custom loss function, ensuring nuanced SCD with heightened accuracy. In addition, Ding et al. [55] introduced the Semantic Change Transformer (SCanFormer), a CSWin Transformer adaptation, to explicitly learn semantic transitions. The model has a triple encoder-decoder architecture to enhance spatial features and incorporates spatio-temporal dynamic and task-specific prior information, leading to a decrease in learning disparities. This approach significantly boosts accuracy on benchmarks, adeptly managing scenarios with sparse labels.
@@
Transformer-based BDA methods are currently understudied. Chen et al. [31] proposed the first pure Transformer-based BDA architecture called DamFormer. It has a Siamese Mix Transformer (MiT) [35] encoder and lightweight all-MLP decoder to efficiently process multi-temporal image pairs. DamFormer employs a multi-temporal fusion module for improved information integration, and a dual-task decoder for precise building localization and damage classification. By adopting strategies like weight sharing and innovative fusion mechanisms, it efficiently handles high-resolution imagery with lower computational costs, making it effective and streamlined for BDA.
@@
Although Transformer-based methods have significantly enhanced CD accuracy through their powerful nonlocal modeling capabilities, they come with substantial computational resource demands. Specifically, as the number of tokens increases, the consumption of computational resources escalates quadratically, posing challenges for processing large-scale input, such as optical high-resolution imagery. Recently, the emerging Mamba architecture not only rivals Transformers in non-local modeling capabilities but also maintains computational overhead at a linear growth rate, which opens up the possibility of efficient and accurate CD.